The accounting records of Wall’s China Shop reflected the following balances as of January 1, Year 3: Cash $ 18,000 Beginning inventory 17,835 (205 @ $87) Common stock 14,500 Retained earnings 21,335 The following five transactions occurred in Year 3: First purchase (cash): 125 units @ $89 Second purchase (cash): 195 units @ $97 Sales (all cash): 360 units @ $191 Paid $13,800 cash for salaries expense Paid cash for income tax at the rate of 25 percent of income before taxes. Required: a. Compute the cost of goods sold and ending inventory, assuming (1) FIFO cost flow, (2) LIFO cost flow, and (3) weighted-average cost flow. Compute the income tax expense for each method. b. Use a vertical model to show the Year 3 income statement, balance sheet, and statement of cash flows under FIFO, LIFO, and weighted average. (Hint: Record the events under an accounting equation before preparing the statements.)
Answer #1
Total inventory available = 47875
Average cost of inventory = 47875/525 = 91.19
Closing inventory (525-360=165 units):
FIFO = 165*97 = 16005
LIFO = 165*87 = 14355
WA = 165 * 91.19 = 15046
COGS:
FIFO = 47875 - 16005 = 31870
LIFO = 47875 - 14355 = 33520
WA = 47875 - 15046 = 32829
Accounting Equations:
FIFO:
Cash inven = CS RE
18000 17835 = 14500 21335
Purchases -11125 11125 =
Purchases -18915 18915 =                                      
Sales 68760                   =                   68760
COGS -31870 = -31870
Salaries paid -13800 = -13800
Income tax paid -5772.5 = -5772.5
Balance 37147.5 16005 = 14500 38652.5
LIFO:
Cash inven = CS RE
18000 17835 = 14500 21335
Purchases -11125 11125 =
Purchases -18915 18915 =                                      
Sales 68760                   =                   68760
COGS -33520 = -33520
Salaries paid -13800 = -13800
Income tax paid -5360 = -5360
Balance 37560 14355 = 14500 37415
WAM:
Cash inven = CS RE
18000 17835 = 14500 21335
Purchases -11125 11125 =
Purchases -18915 18915 =                                      
Sales 68760                   =                   68760
COGS -32829 = -32829
Salaries paid -13800 = -13800
Income tax paid -5532.75 = -5532.75
Balance 37387.25 15046 = 14500 37933.25
Income Statement:
FIFO LIFO WAM
Sales 68760 68760 68760
Less:COGS -31870 -33520 -32829
Gross profit 36890 35240 35931
Salaries paid -13800 -13800 -13800
Income before tax 23090 21440 22131
Less: Tax 5772.5 5360 5532.75
Net Income 17317.5 16080 16598.25
Balance Sheet:
Assets: FIFO LIFO WAM
Cash 37147.5 37560 37387.25
Inventory 16005 14355 15046
Total Assets 53152.5 51915 52433.25
Liabilities:
Com Stock 14500 14500 14500
Ret. Earnings 38652.5 37415 37933.25
total Liabilities 53152.5 51915 52433.25
Cash Flow Statement:
FIFO LIFO WAM
Cash Op Balance 18000 18000 18000
Purchases -11125 -11125 -11125
Purchases -18915 -18915 -18915
Sales 68760 68760 68760
Salaries paid -13800 -13800 -13800
Income tax paid -5772.5 -5360 -5532.75
Cash Cl Balance 37147.5 37560 37387.25